### v4

# LUDVIG: Learning-free Uplifting of 2D Visual features to Gaussian Splatting scenes

Abstract
We address the problem of extending the capabilities of vision foundation models such as DINO, SAM, and CLIP, to 3D tasks. Specifically, we introduce a novel method to uplift 2D image features into Gaussian Splatting representations of 3D scenes. Unlike traditional approaches that rely on minimizing a reconstruction loss, our method employs a simpler and more efficient feature aggregation technique, augmented by a graph diffusion mechanism. Graph diffusion refines 3D features, such as coarse segmentation masks, by leveraging 3D geometry and pairwise similarities induced by DINOv2. Our approach achieves performance comparable to the state of the art on multiple downstream tasks while delivering significant speed-ups. Notably, we obtain competitive segmentation results using generic DINOv2 features, despite DINOv2 not being trained on millions of annotated segmentation masks like SAM. When applied to CLIP features, our method demonstrates strong performance in open-vocabulary object localization tasks, highlighting the versatility of our approach. [^1] Project page: https://juliettemarrie.github.io/ludvig
1 Introduction
The field of image understanding has recently seen remarkable progress, driven by large pretrained models such as CLIP (Radford et al., 2021), DINO (Caron et al., 2021; Oquab et al., 2024), or SAM (Kirillov et al., 2023). A key factor behind their exceptional generalization capabilities lies in the vast size of their training datasets, often composed of millions or even billions of samples.
3D scene representation has also advanced through machine learning approaches like NeRF (Mildenhall et al., 2021) and model fitting techniques such as Gaussian Splatting (Kerbl et al., 2023). These methods typically rely on a few dozen views of the scene captured from different angles. While the resulting reconstructions effectively capture both appearance and geometrical information, they are not directly applicable to semantic tasks, which has led to further developments.
The complementarity of these two families of approaches has indeed recently been exploited by numerous methods that integrate geometry and semantics by uplifting image-level features extracted by large pretrained models into 3D NeRF or Gaussian Splatting representations. This has led to a surge in methods for tasks such as language-guided object retrieval (Kerr et al., 2023; Liu et al., 2023; Zuo et al., 2024), scene editing (Kobayashi et al., 2022; Chen et al., 2024; Fan et al., 2023), or semantic segmentation (Cen et al., 2023c; Ye et al., 2024a; Ying et al., 2024).
The main limitation of most previous approaches lies in their dependence on optimization, which requires an iterative process to learn a scene-specific 3D representation by minimizing a reprojection error across all training views. While this loss function is intuitive, a faster and more straightforward method for transferring 2D generic visual features to already trained Gaussian splatting 3D models would be preferable, which is the purpose of this work.
In this paper, we demonstrate that a simple, learning-free process is highly effective for uplifting 2D features or semantic masks into 3D Gaussian Splatting scenes. This process, which can be viewed as an ‘inverse rendering’ operation, is both computationally efficient and adaptable to any feature type. We showcase its effectiveness by uplifting visual features from DINOv2 (Oquab et al., 2024; Darcet et al., 2024), semantic masks from SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024), and language features from CLIP (Ilharco et al., 2021). Then, we show that a graph diffusion mechanism (Kondor & Lafferty, 2002; Smola & Kondor, 2003) is helpful for feature refinement in 3D scenes. This mechanism is rooted in spectral graph theory and used in spectral clustering techniques (Belkin & Niyogi, 2001; Shi & Malik, 2000; Meila & Shi, 2000). In the context of our work, it transforms coarse segmentation inputs, such as scribbles or alignment scores between visual features and a text query, into accurate 3D segmentation masks without the need for segmentation models such as SAM. When evaluated on segmentation and open-vocabulary object localization, our method achieves results comparable to state-of-the-art techniques while being significantly faster than previous approaches relying on optimization.
To summarize, our contributions are threefold: (i) we introduce a simple, learning-free uplifting approach that can be directly integrated into the rendering process, achieving state-of-the-art results when applied to SAM-generated semantic masks (Sec.4.1). (ii) we demonstrate that using graph diffusion based on uplifted DINOv2 features yields competitive results for foreground/background and open-vocabulary object segmentation, despite DINOv2 not being explicitly trained for segmentation like SAM (Secs. 4.14.2). (iii) we show that combining SAM with graph diffusion achieves state-of-the-art results on open-vocabulary object segmentation tasks (Sec.4.2).
2 Related work
Learning 3D semantic scene representations with NeRF.
NeRF (Mildenhall et al., 2021) uses a multilayer perceptron to predict the volume density and radiance for any given 3D position and viewing direction. Such representation can naturally be extended to semantic features. The early works N3F (Tschernezki et al., 2022) and DFF (Kobayashi et al., 2022) distill DINO 2D (i.e., image-level) features (Caron et al., 2021) in scene-specific NeRF representations. Kobayashi et al. (2022) also distill LSeg (Li et al., 2022) a CLIP-inspired language-driven model for semantic segmentation. Shortly after, LERF (Kerr et al., 2023) and 3D-OVS (Liu et al., 2023) learned 3D CLIP (Radford et al., 2021) and DINO (Caron et al., 2021) features jointly for open-vocabulary segmentation. These works were extended to other pretrained models such as latent diffusion models (Ye et al., 2023) or SAM (Kirillov et al., 2023) for semantic segmentation (Cen et al., 2023c; Ying et al., 2024).
Learning 3D semantic scene representations with Gaussian splatting.
Subsequent work has relied on the more recent Gaussian splatting method (Kerbl et al., 2023), achieving high-quality novel-view synthesis while being orders of magnitude faster that NeRF-based models. Several tasks have been addressed such as semantic segmentation using SAM (Cen et al., 2023b; Ye et al., 2024a; Kim et al., 2024), language-driven retrieval or editing using CLIP combined with DINO (Zuo et al., 2024) or SAM (Ye et al., 2023), scene editing using diffusion models (Chen et al., 2024; Wang et al., 2024), and 3D-aware finetuning (Yue et al., 2024). These works learn 3D semantic representations by minimizing a reprojection loss. As a single scene can be represented by over a million Gaussians, such optimization-based techniques have strong memory and computational limitations. To handle these, FMGS (Zuo et al., 2024) employs a multi-resolution hash embedding (MHE) of the scene for uplifting DINO and CLIP representations, Feature 3DGS (Zhou et al., 2024) learns a $1\times 1$ convolutional upsampler of Gaussians’ features distilled from LSeg and SAM’s encoder, and LangSplat (Qin et al., 2024) learns an autoencoder to reduce CLIP feature dimension from 512 to 3. In contrast, our approach requires no learning, which significantly speeds up the uplifting process and reduces the memory requirements.
Direct uplifting of 2D features into 3D.
Direct uplifting from 2D to 3D has been explored in prior works. OpenGaussian (Wu et al., 2024) proposes an instance-level 3D-2D feature association method that links 3D points to 2D masks, while GaussianEditor (Chen et al., 2024) uplifts 2D SAM masks into 3D to selectively optimize Gaussians for editing tasks (see details in Appendix Sec.C.3). Both approaches are specific to segmentation and are ill-suited for uplifting generic representations. Semantic Gaussians (Guo et al., 2024) pairs 2D pixels with 3D Gaussians along each pixel’s ray based on depth information but relies on a learned 3D convolutional network. In contrast, our approach is both applicable to any 2D representation and entirely parameter-free.
Leveraging 3D information to better segment in 2D.
Most prior works focusing on semantic segmentation leverage 2D models specialized for this task. The early work of Yen-Chen et al. (2022) uplifts learned 2D image inpainters by optimizing view consistency over depth and appearance. Subsequent works have mostly relied on uplifting either features from SAM’s encoder (Zhou et al., 2024), binary SAM masks (Cen et al., 2023c; b), or SAM masks automatically generated for all objects in the image (Ye et al., 2024a; Ying et al., 2024; Kim et al., 2024). The latter approach is computationally expensive, as it requires querying SAM on a grid of points over the image. It also requires matching inconsistent mask predictions across views, with e.g. a temporal propagation model (Ye et al., 2024a) or a hierarchical learning approach (Kim et al., 2024), which introduces additional computational overhead. In this work, we focus on single instance segmentation and show that our uplifted features are on par with the state of the art (Cen et al., 2023c; b; Ying et al., 2024). Standing out from prior work uplifting DINO features (Tschernezki et al., 2022; Kobayashi et al., 2022; Goel et al., 2023; Kerr et al., 2023; Liu et al., 2023; Ye et al., 2023; Zuo et al., 2024), we quantitatively show that DINOv2 features can be used on their own for semantic segmentation and rival SAM-based models through a simple graph diffusion process that leverages 3D geometry.
Learning 3D CLIP features for open-vocabulary object localization.
For learning 3D CLIP features, prior works also leverage vision models such as DINO or SAM. DINO is used to regularize and refine CLIP features (Kerr et al., 2023; Liu et al., 2023; Zuo et al., 2024; Shi et al., 2024), while SAM is employed for generating instance-level CLIP representations (Qin et al., 2024). These approaches suffer from high computational costs, resorting to either dimensionality reduction or efficient multi-resolution embedding representations, and usually run for a total of one to two hours for feature map generation and 3D feature optimization. In contrast, our approach bypasses the high computational cost of gradient-based optimization and, combined with graph diffusion, is an order of magnitude faster than these prior works.
3 Uplifting 2D visual representations into 3D
In this section, we present a simple yet effective method for lifting 2D visual features into 3D using Gaussian splatting and discuss its relation with more expensive optimization-based techniques.
3.1 Background on Gaussian splatting
Scene representation.
The Gaussian splatting method consists in modeling a 3D scene as a set of $n$ Gaussians densities ${\mathcal{N}}_{i}$, each defined by a mean $\mu_{i}$ in $\mathbb{R}^{3}$, a covariance $\Sigma_{i}$ in $\mathbb{R}^{3\times 3}$, an opacity $\sigma_{i}$ in $(0,1)$, and a color function $c_{i}({d})$ that depends on the viewing direction $d$ [^2] In this work, the viewing direction $d$ refers to the full camera pose, defined by its extrinsic and intrinsic parameters..
A 2D frame at a given view is an image $\hat{I}_{d}$ rendered by projecting the 3D Gaussians onto a 2D plane, parametrized by the viewing direction $d$. This projection accounts for the opacity of the Gaussians and the order in which rays associated with each pixel pass through the densities. More precisely, a pixel $p$ for a view $d$ is associated to an ordered set ${\mathcal{S}}_{{d},p}$ of Gaussians and its value is obtained by their weighted contributions:
$\hat{I}_{{d}}(p)=\sum_{i\in{\mathcal{S}}_{d,p}}c_{i}(d)w_{i}(d,p).$ (1)
The above weights are obtained by $\alpha$-blending, i.e. $w_{i}(d,p)=\alpha_{i}(d,p)\prod_{j\in\mathcal{S}_{d,p},j<i}\left(1-\alpha_{j}(d,p)\right)$, where the Gaussian contributions $\alpha_{i}(d,p)$ are computed by multiplying the opacity $\sigma_{i}$ by the Gaussian density ${\mathcal{N}}_{i}$ projected onto the 2D plane at pixel position $p$.
Scene optimization.
Let $I_{1},\dots,I_{m}$ be a set of 2D frames from a 3D scene and $d_{1},\dots,d_{m}$ the corresponding viewing directions. Gaussian Splatting optimizes the parameters involved in the scene rendering function described in the previous section. This includes the means and covariances of the Gaussian densities, their opacities, and the color function parametrized by spherical harmonics. Denoting by $\theta$ these parameters, the following reconstruction loss is used
$\min_{\theta}\frac{1}{m}\sum_{k=1}^{m}\mathcal{L}(I_{k},\hat{I}_{d_{k},\theta}),$ (2)
where $\hat{I}_{d_{k},\theta}$ is the rendered frame of the scene in the direction $d_{k}$, as in Eq. (1), by using the parameters $\theta$, and $\mathcal{L}$ is a combination of $\ell_{1}$ and SSIM loss functions (Kerbl et al., 2023).
3.2 Uplifting 2D feature maps into 3D
Given a set of $m$ 2D training frames and the corresponding 3D scene obtained by Gaussian Splatting, our goal is to compute generic features $f_{i}$ in $\mathbb{R}^{c}$ for each Gaussian $i$, which would be effective for solving future downstream tasks, e.g., high-resolution semantic segmentation for new frames of the scene, or open-vocabulary object localization. In other words, $f_{i}$ can be seen as an extension of the color function $c_{i}$, even though, for simplicity, we do not consider view-dependent features in this work.
A natural approach is to consider a pretrained vision model that provides 2D feature maps for each of the $m$ frames used in Gaussian Splatting, and then devise a technique to uplift these 2D feature maps into 3D. This uplifting principle can also be directly applied to semantic masks instead of generic features, as demonstrated in Section 5. Interestingly, once the features $f_{i}$ are computed for each Gaussian $i$, it is possible to render two-dimensional feature maps for any new view, at a resolution that can be much higher than the feature maps computed for the $m$ training frames.
Uplifting with simple aggregation.
We construct uplifted features for each 3D Gaussian of the 3D Gaussian Splatting scene as a weighted average of 2D features from all frames. Each 2D feature $F_{d,p}$ from a frame at a given viewing direction $d$ and pixel $p$ contributes to the feature $f_{i}$ by a factor proportional to the rendering weight $w_{i}(d,p)$, if the Gaussian $i$ belongs to the ordered set $\mathcal{S}_{d,p}$ associated to the view/pixel pair $(d,p)$. Denoting $S_{i}=\{(d,p),i\in\mathcal{S}_{d,p}\}$ the set of view/pixel pairs contributing to the feature $f_{i}$, the resulting features are defined as follows:
$f_{i}=\sum_{(d,p)\in\mathcal{S}_{i}}\bar{w}_{i}(d,p)F_{d,p}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{with}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \bar{w}_{i}(d,p)=\frac{w_{i}(d,p)}{\sum_{(d,p)\in\mathcal{S}_{i}}w_{i}(d,p)},$ (3)
We can interpret this equation as a normalized version of the transposed rendering operation over the $m$ viewing directions. More precisely, the rendering of any view-independent collection of features ${\bf{f}}=(f_{i})$ attached to the $n$ Gaussians into the $m$ training frames can be represented as a linear operator $W$ acting on the collection $\bf{f}$ and returning a collection of 2D feature maps ${\bf{\hat{F}}}=(\hat{F}_{d,p})$, see Eq. (4) below. Here, non-zero entries of the matrix $W$ consists of all rendering weights $w_{i}(d,p)$ when $(d,p)\in\mathcal{S}_{i}$ is placed at row $(d,p)$ and column $i$, and $\mathbf{\hat{F}}$ is a 2D matrix containing all (flattened) 2D feature maps generated for all cameras poses, with $\mathbf{\hat{F}}_{d,p}$ the feature of pixel $p$ from view at direction $d$. Similarly, the uplifting expression introduced in Eq. (3) can be expressed in terms of the transpose of $W$ and a diagonal matrix $D$ of size $m$ representing the normalization factor and whose diagonal elements are obtained by summing over the rows of $W$ as in Eq. (5) below:
Rendering to $m$ frames
${\bf\hat{F}}=W{\bf{f}},$ (4)
Uplifting from $m$ frames
${\bf f}=D^{-1}W^{\top}{\bf{F}}.$ (5)
It is important to note that $W$ and $D$ are not explicitly constructed. Instead, they are computed by calling the forward rendering function for Gaussian Splatting and replacing the color vectors by the feature vectors. All these operations are performed within the CUDA rendering process and are illustrated in Figure 1.
Figure 1: Illustration of the inverse and forward rendering. In the inverse rendering (or uplifting) phase, features ${\bf{f}}$ are created for each 3D Gaussian by aggregating coarse 2D features ${\bf F}$ over all viewing directions. For forward rendering, the 3D features ${\bf{f}}$ are projected on any given viewing direction as in regular Gaussian Splatting. The rendering weight $\bar{w}_{i}(d,p)$ represents the relative influences of the Gaussian $i$ on pixel $p$, defined in Eq. (3).
Connection with optimization-based inverse rendering.
An alternative approach to uplifting 2D features ${\bf F}$ is to minimize a reconstruction objective $\mathcal{L}({\bf f})$, where the goal is to find uplifted features ${\bf f}$ whose rendering closely matches the original 2D features ${\bf F}$ (Tschernezki et al., 2022; Kerr et al., 2023; Zuo et al., 2024). A natural choice is to minimize the mean squared error between the 2D features ${\bf F}$ and the rendered ones ${\bf\hat{F}}$ as defined by Eq. (4):
$\min_{{\bf f}}\mathcal{L}({\bf f}):=\frac{1}{2}\|{\bf F}-W{\bf f}\|^{2}.$ (6)
Such an approach requires an optimization procedure which is costly compared to our proposed uplifting method. Nevertheless, it is possible to interpret the proposed uplifting scheme in Eq. (5) as a single pre-conditioned gradient descent step on the reconstruction objective, starting from a ${\bf 0}$ feature, i.e., ${\bf f}=-D^{-1}\nabla\mathcal{L}({\bf 0})$. In practice, we found that performing more iterations on the objective $\mathcal{L}({\bf f})$ did not improve the quality of the features, thus suggesting that the computationally cheaper scheme in Eq. (5) is already an effective approach to uplifting.
Gaussian filtering
The normalization term $\beta_{i}=\sum_{d,p\in\mathcal{S}_{i}}w_{i}(d,p)$ serves as an estimator of the relative importance of each Gaussian in the scene. Therefore, it can be used as a criterion to prune the set of Gaussians for memory efficiency. In our experiments, we filter out half of the Gaussians based on $\beta_{i}$ and observe no qualitative nor quantitative degradation of the results. This approach is inspired by prior work on efficient Gaussian Splatting representation such as proposed by Fan et al. (2023) that also prunes Gaussians based on their contribution to each pixel in the training frames.
3.3 Enriching features by graph diffusion
DINOv2 features have shown remarkable performance on semantic segmentation tasks with simple linear probing (Oquab et al., 2024), making them a good candidate to enrich features that lack such a property like CLIP (Wysoczańska et al., 2024; Zuo et al., 2024; Liu et al., 2023). Inspired by spectral clustering techniques (Shi & Malik, 2000; Kondor & Lafferty, 2002; Belkin & Niyogi, 2001), we propose to diffuse features that have been uplifted to 3D. This process aims to align semantic features with the scene layout and object boundaries implied by DINOv2. Contrary to the aforementioned methods, we are the first to perform this landscaping with DINOv2 directly in the 3D scene, hereby taking 3D geometry into account as well.
Graph construction
We construct a graph from uplifted DINOv2 features $f$ in $\mathbb{R}^{n}$. Nodes are given by the 3D Gaussians. Edges, represented by a matrix $A$ of size $n\times n$, encode both the 3D Euclidean geometry between the nodes and the similarity between their DINOv2 features. More precisely, we extract the $k$ nearest neighbors $\mathcal{N}(i)$ for each node $i$, as measured by the Euclidean distance between the centers of the 3D Gaussians. Two nodes $i$ and $j$ in the graph are linked by an edge if $i\in\mathcal{N}(j)$ or $j\in\mathcal{N}(i)$, and the edge is assigned the following weight:
$A_{ij}=S_{f}(f_{i},f_{j})\,P(f_{i})^{\frac{1}{2}}P(f_{j})^{\frac{1}{2}},$ (7)
with $S_{f}(f_{i},f_{j})$ a local similarity between features $f_{i}$ and $f_{j}$, typically defined as a RBF kernel. For tasks requiring diffusion to be confined to a specific object instance, we prevent leakage into the background by introducing a node-wise unary regularization term $P(f_{i})$ which quantifies the similarity between the node feature $f_{i}$ and the features of the object of interest. Details on $S_{f}$ and $P$ are provided in Appendix  A.3.
Diffusion on the graph.
Given initial 3D features $g_{0}$ in $\mathbb{R}^{n}$, which we aim to improve by using information encoded in $A$ (3D geometry and DINOv2 similarities), we perform $T$ diffusion steps to construct a sequence of diffused features $(g_{t})_{1\leq t\leq T}$ defined as follows:
$g_{t+1}=A\tilde{g}_{t},\quad\tilde{g}_{t}={g_{t}}/{\|g_{t}\|_{2}},$ (8)
This can be seen as performing a few steps of the power method, making $g_{0}$ closer to the dominant eigenspace of $A$. Note that depending on the downstream task, $g_{0}$ may represent generic features or task-specific features such as coarse 3D segmentation masks.
4 From 3D Uplifting to downstream tasks
In this section, we describe our approach for extracting and uplifting features from SAM, DINOv2, and CLIP, as well as evaluating these features on two downstream tasks: segmentation and open-vocabulary object localization. As in Sec. 3, we are given a set of 2D frames $I_{1},\dots,I_{m}$, with viewing directions $d_{1},\dots,d_{m}$, and a 3D Gaussian Splatting representation of the scene.
4.1 Multi-view Segmentation
We assume that a foreground mask of the object to be segmented is provided on the reference frame $I_{1}$. The foreground masks are either scribbles or a whole reference mask of the object, both of which define a set of foreground pixels $\mathcal{P}$. In the following, we present the proposed approaches for segmentation using SAM and DINOv2 features, based on both types of foreground masks.
Multi-view segmentation with SAM.
SAM (Kirillov et al., 2023; Ravi et al., 2024) is a powerful model that can generate object segmentation masks from point prompts, on a single 2D image. Aggregating SAM 2D segmentation masks in 3D allows for cross-view consistency and improves single-view segmentation results. We proceed by generating 2D feature maps based on SAM segmentation masks of each training frame while only relying on the foreground mask for the reference frame $I_{1}$. The 2D feature maps are generated by constructing several sets of point prompts on each training frame which are then provided to SAM to obtain several segmentation masks. The point prompts are obtained using the foreground mask provided on the reference frame as described in Appendix A.1. Averaging the resulting segementation masks for each frame results in the final 2D SAM feature maps. These are then uplifted using the aggregation scheme in Sec. 3.2. Our final prediction is obtained by rendering the uplifted feature maps into the target frame and thresholding.
Multi-view segmentation with DINOv2.
We construct 2D feature maps at the patch level using DINOv2 with registers (Darcet et al., 2024) and uplift them into a high resolution and fine-grained 3D semantic representation which is then used for segmentation. The 2D feature maps are constructed using a combination of a sliding windows mechanism and dimensionality reduction of the original DINOv2 features as described in Appendix A.2 and illustrated in Figure 4 therein. This approach enhances the granularity of spatial representations by aggregating patch-level representations to form pixel-level features. To favor the first principal components, known to focus on the foreground objects (Oquab et al., 2024), the features are re-weighted by the eigenvalues of the PCA decomposition. The 2D feature maps from the $m$ training views are uplifted using Eq. (3) and the resulting 3D features are then re-projected into any viewing direction $d$ using Eq. (4) to compute rendered 2D features $(\hat{F}_{d,p})$. To obtain segmentation masks, we define a predictor score $P(\hat{F}_{d,p})$ as the likelihood that a 2D pixel $p$ belongs to the foreground, based on its feature $\hat{F}_{d,p}$. The score $P$ is obtained by comparing the rendered features $(\hat{F}_{d,p})$ with foreground features $\mathcal{F}_{\text{ref}}:=(\hat{F}_{d_{1},p})_{p\in\mathcal{P}}$ corresponding to the foreground mask from the reference frame $I_{1}$, and the final segmentation mask is then obtained by thresholding. More details are provided in Appendix A.2.
Enhancing segmentation with DINOv2 using 3D graph diffusion.
DINOv2 provides generic visual features that do not explicitly include segmentation information, unlike models such as SAM that were specifically trained for such a task. Consequently, 2D projections of uplifted DINOv2 features might fail to separate distinct objects that have similar features. This challenge can be mitigated by incorporating 3D spatial information.
To this end, we propose to leverage the graph diffusion process introduced in Section 3.3 and illustrated in Figure 3. We set the initial vector of weights $g_{0}\in\mathbb{R}^{n}$ of the graph diffusion algorithm to be a coarse estimation of the contribution of each Gaussian to the final segmentation mask. This initial weight vector is computed by uplifting the 2D foreground mask (either scribbles or a reference mask) from the reference frame into 3D using Eq. (3), normalizing and thresholding them (see Appendix A.3) The nodes for which $g_{0}$ has a positive value define a set of anchor nodes $\mathcal{M}$ that are more likely to contribute to the foreground. The regularization term $P$ appearing in Eq. (7) is obtained by comparing the uplifted features with anchor features obtained using the foreground mask as described in the appendix. For this task, we binarize $A$ with a fixed threshold (set to $10^{-5}$). After the $T$ diffusion steps, we recover the nodes $\mathcal{S}$ in $g_{T}$ with strictly positive values (i.e., those reachable after $T$ iterations). The final weight is defined as $h_{i}=P(f_{i})$ if $i\in\mathcal{S}$ and 00 otherwise. Segmentation is then performed by projecting $\mathbf{h}=(h_{i})$ into 2D and thresholding.
4.2 Open-vocabulary object detection
Following Kerr et al. (2023); Qin et al. (2024); Zuo et al. (2024), we uplift CLIP features (Ilharco et al., 2021) for open-vocabulary object localization. CLIP effectively aligns images and text in a shared representation space. As a measure of alignment, we use the relevancy score introduced by LERF (Kerr et al., 2023), which quantifies the similarity between a CLIP visual feature and a text query.
Construction of CLIP feature maps.
We follow common practice (Kerr et al., 2023; Zuo et al., 2024) and construct multi-resolution CLIP 2D feature maps by querying CLIP on a grid of overlapping patches at different scales and aggregating the resulting representations. As in Zuo et al. (2024), rather than keeping the different representations separate, we aggregate them with a simple average pooling. These multi-resolution CLIP features are uplifted into 3D using Eq. (3).
Relevancy scores.
After uplifting CLIP features, we compute relevancy scores for each Gaussian’s feature to text queries embedded by CLIP. These relevancy scores can then be projected into 2D and used for both localization and segmentation. For localization, we choose the pixel with the highest relevancy score. For segmentation, we predict a SAM mask by selecting point prompts among pixels with the highest relevancy score. Specifics on the computation of relevancy scores and segmentation masks are provided in Appendix A.4.
Refining relevancy scores with DINOv2 graph diffusion.
We refine 3D relevancy scores with the diffusion process described in Sec. 3.3. To this end, DINOv2 features are also uplifted, and the similarity matrix is built as in Eq. (7), with the unary term $P$ constructed using a logistic regression over thresholded relevancies, see details in Appendix A.4. The diffusion process propagates CLIP relevancies to Gaussians with similar DINOv2 features. The resulting 3D relevancy scores span the object of interest without covering other objects with similar features and show a strong decay at the object’s borders, as defined by DINOv2’s feature landscape, resulting in improved segmentation results.
5 Experiments
5.1 Experiment details
3D scene training and pruning.
All scenes are trained using the original Gaussian Spatting implementation (Kerbl et al., 2023) with default hyperparameters. For memory efficiency, half of the Gaussians are filtered out based on their importance, as described in Sec. 3.2.
2D vision models.
Our experiments are conducted using DINOv2’s ViT-g with registers (Darcet et al., 2024), SAM (Kirillov et al., 2023), SAM 2 (Ravi et al., 2024) and the OpenCLIP ViT-B/16 model (Ilharco et al., 2021).
Segmentation tasks.
We consider two segmentation tasks: i) Neural Volumetric Object Selection (NVOS,  Ren et al. 2022), which is derived from the LLFF dataset (Mildenhall et al., 2019), and ii) SPIn-NeRF, which contains a subsets of scenes from NeRF-related datasets (Knapitsch et al., 2017; Mildenhall et al., 2019; 2021; Yen-Chen et al., 2022; Fridovich-Keil et al., 2022). The NVOS dataset consists of forward-facing sequences where the task is to predict the segmentation mask of a labeled frame based on reference scribbles from another frame. SPIn-NeRF contains both forward-facing and 360-degree scenes, in which all frames are labeled with segmentation masks, and the standard evaluation protocol uses the segmentation mask from the first frame as reference to label the subsequent frames.
Open-vocabulary object localization
We evaluate our approach on the LERF dataset (Kerr et al., 2023), which includes localization and segmentation tasks on complex in-the-wild scenes. We report our results on the extended evaluation task introduced by LangSplat (Qin et al., 2024) containing additional challenging localization samples, and also follow their evaluation protocol.
Evaluation and hyperparameter tuning.
Results are averaged over 3 independent runs, and hyperparameters are selected as follows. Segmentation with 3D SAM masks requires setting a threshold for foreground/background pixel assignment, and optionally choosing one of the three masks proposed by SAM (representing different possible segmentations of the object of interest). Segmentation with DINOv2 relies on two parameters for defining $S_{f}$ and $P$ in graph diffusion, and on the threshold for foreground/background pixel assignment. These parameters are automatically chosen as described in Appendix Secs. A.1 and A.3.
For the open-vocabulary object localization tasks, all parameters including those used for graph diffusion and thresholds for segmentation are either fixed for all scenes or automatically selected, as detailed in Appendix A.4.
(a) RGB image
(b) Single-view DINOv2 PCA
(c) Multi-view DINOv2 PCA
Figure 2: PCA visualizations. The DINOv2 patch-level representations (middle) predicted from the RGB images (left) are aggregated into highly detailed 3D representations (right) using Eq. (3).
(a) RGB image
(b) 3 steps
(c) 7 steps
(d) 100 steps
Figure 3: Illustration of the diffusion process. 2D projection of the weight vector $g_{t}$ (white) and unary regularization term (red) at different diffusion steps $t$. The diffusion process allows filtering out unwanted objects that have similar features to the object of interest (such as the two smaller skulls on horns, bottom-row), but are disconnected in space. The regularization term (red background) prevents leakage from the object to the rest of the scene (such as through the fern’s trunk, top-row).
5.2 Qualitative results
DINOv2 feature uplifting.
First, we illustrate the effectiveness of our simple uplifting approach. Figure 2 shows the first three PCA components (one channel per component) over DINOv2’s patch embeddings. The coarse patch-level representations from every view (middle) are aggregated using Eq. (5) to form a highly detailed 3D semantic representation, and reprojected into 2D (right) using Eq. (4). The aggregation is very fast, being directly implemented in the Gaussian Splatting CUDA-based rendering process, and takes about 1.3ms per view and feature dimension. The first principal component (encoded in the red channel) mostly captures the foreground object, and the subsequent ones allow refining the foreground representations and delivering a detailed background. In the appendix, we provide additional comparative visualizations of our learned 3D features (Figure 8) and of 3D segmentation for scene editing (Figure 7).
Graph diffusion.
Figure 3 illustrates the effectiveness of the diffusion process. The graph nodes are initialized with the reference scribbles, and the diffusion spreads through the object of interest, stopping at its borders. As illustrated with the case of Horns, diffusion filters out unwanted objects that are similar to the object of interest (here, the two skulls on the side). In the Fern scene, diffusion progressively spreads through the branches to their extremities, with the regularization (red background) constraining it to the trunk and preventing leakage, even after a large number of iterations. Appendix Figure 6 also illustrates this for the Flower and Trex scenes: diffusion rapidly spreads, achieving near-full coverage after only 5 steps before reaching all the much smaller Gaussians on the border, allowing for a refined segmentation.
NVOS MVSeg SA3D-TRF SA3D-GS SAGA OmniSeg3D LUDVIG (Ours)
3D representation: TensoRF GS GS NeRF GS
Uplifting: SAM SAM SAM SAM DINOv2 SAM SAM2
NVOS 70.1 - 90.3 92.2 92.6 91.7 92.4 91.3 91.3
SPIn-NeRF - 90.9 93.7 93.2 93.4 94.3 93.8 93.8 93.8
Table 1: Segmentation (IoU) on NVOS (Ren et al., 2022) and SPIn-NeRF (Mirzaei et al., 2023).
Geometry only Single view Uplifting Uplifting + Graph diffusion
Reference mask DINOv2 SAM2 DINOv2 SAM2 DINOv2
73.1 88.5 88.6 91.6 93.8 93.8
Table 2: Segmentation (IoU) on SPIn-NeRF (Mirzaei et al., 2023). We compare purely geometrical reference mask uplifting and reprojection and single-view prediction, feature/mask uplifting or graph diffusion leveraging DINOv2 or SAM2.
5.3 Segmentation results
In this section, we quantitatively evaluate the segmentation task on NVOS (Ren et al., 2022) and SPin-NeRF (Mirzaei et al., 2023). We evaluate segmentation based on SAM and SAM2 mask uplifting, and on DINOv2 feature uplifting combined with graph diffusion. We compare our segmentation results to the current state of the art: SA3D (Cen et al., 2023c), SA3D-GS (Cen et al., 2023b), SAGA (Cen et al., 2023a), OmniSeg3D (Ying et al., 2024). All these methods are specifically designed for uplifting the 2D segmentation masks produced by SAM into 3D using gradient-based optimization of a projection loss. We also report results from NVOS (Ren et al., 2022) and MVSeg (Yen-Chen et al., 2022), who respectively introduced the NVOS and SPIn-NeRF datasets.
Results.
Table 1 reports the average IoU across all scenes for NVOS and SPIn-NeRF. Per-scene results can be found in Appendix Tables 6 and  7. Our results are comparable to the state-of-the-art, while not relying on gradient-based optimization. Surprisingly, our segmentation with DINOv2 using graph diffusion also gives results on par with models leveraging SAM masks. Compared to SAM, DINOv2 better captures complex objects, but sometimes also captures some background noise. This can be seen in Appendix Figure 5 with the example of the T-Rex: while SAM misses out the end of the tail as well as the end of the ribs, DINOv2 captures the whole Trex, but also captures part of the stairs behind. Our lower segmentation results compared to OmniSeg’s can partly be attributed to the poor Gaussian Splatting reconstruction of highly specular scenes, such as the Fork. As also noted by Cen et al. (2023a), the reconstruction includes semi-transparent Gaussians floating over the object, attempting to represent reflections or surface effects, which are challenging to capture using standard rasterization techniques (Jiang et al., 2024).
Ablation study.
We compare our segmentation protocol using DINOv2 and SAM2 to multiple simpler variants. More precisely, we evaluate i) a purely geometrical variant that reprojects the reference mask on the other views, without using SAM2 or DINOv2, ii) single-view segmentation in 2D based on SAM2 or DINOv2 2D predictions, iii) uplifting DINOv2 features or SAM2 masks into 3D then rendering them for segmentation, and iv) segmenting using graph diffusion over DINOv2 3D feature similarities. Results are reported in Table 2, and per-scene IoU as well as a detailed analysis can be found in Appendix Table 8 and Sec. B.2. We observe that the purely geometrical approach works well on the forward-facing scenes and fails on 360-degree scenes. The single-view variant performs reasonably well on average but, the low resolution of patch-level representations (illustrated in Figure 2) lead to a coarser segmentation. 3D uplifting considerably boosts results compared to single-view approaches, and introducing 3D spatial information through 3D graph diffusion further enhances results on the more challenging 360-degree scenes.
5.4 Open-vocabulary object detection
Table 3 presents results on the open-vocabulary LERF localization and segmentation tasks (Kerr et al., 2023), evaluated on the extended and more challenging version of the dataset introduced by LangSplat (Qin et al., 2024). The reported average running times include feature map generation and 3D feature training whenever relevant. LUDVIG outperforms prior works while being significantly faster than all methods (around 10 times faster). Table 3 reports our results for segmentation using graph diffusion for refining 3D CLIP relevancy scores, followed by segmentation with SAM. Table 4 also reports results with and without using graph diffusion and/or SAM.
A more thorough analysis on running times can be found in Appendix B.1. Additionally, Appendix C.4 provides illustrations of the impact of the 3D diffusion process and 2D SAM segmentation (Figure 10), and comparative visualizations of localization heatmaps with LangSplat and LERF (Figure 11).
Localization (Accuracy) Segmentation (IoU) Time
ramen figurines teatime waldo overall ramen figurines teatime waldo overall (mins)
LERF 62.0 75.0 84.8 72.7 73.6 28.2 38.6 45.0 37.9 37.4 45
LangSplat 73.2 80.4 88.1 95.5 84.3 51.2 44.7 65.1 44.5 51.4 105
LUDVIG 78.9 80.4 94.9 90.9 86.3 58.1 63.3 77.1 58.5 64.3 10
Table 3: LERF Localization and Segmentation. We evaluate on the more challenging dataset introduced by LangSplat (Qin et al., 2024).
SAM Graph diffusion ramen figurines teatime waldo overall
27.8 37.8 38.2 30.4 33.6
42.3 58.0 58.6 42.9 50.4
52.2 51.8 68.9 56.4 57.3
58.1 63.3 77.1 58.5 64.3
Table 4: LERF Segmentation. Results (IoU) with and without using 3D graph diffusion and/or 2D SAM segmentation, evaluated on the dataset introduced by LangSplat (Qin et al., 2024).
6 Concluding remarks and limitations
In this work, we introduce a simple yet effective aggregation mechanism for transferring 2D visual representations into 3D, bypassing traditional optimization-based approach. The aggregation builds upon already trained Gaussian Splatting representations and is implemented within the CUDA rendering process, making 2D-to-3D uplifting as fast as 3D-to-2D rendering. Note however that the quality of 3D features is bound by that of the 3D scene reconstruction. Reconstruction by Gaussian Splatting is notoriously challenging in cases of high specularity (Jiang et al., 2024; Yang et al., 2024), blurred images (Zhao et al., 2024; Lee et al., 2024) or high-frequency regions (Ye et al., 2024b; Zhang et al., 2024). In such scenarios, learning 3D features along with 3D Gaussian Splatting reconstruction may improve scene geometry, opening promising perspectives for future work.
After feature uplifting, our graph diffusion process allows injecting the rich DINOv2 representations to transform coarse 2D or 3D segmentation signals, such as scribbles or CLIP relevancy scores, into accurate 3D segmentation masks. Our CLIP relevancy refinement builds upon prior works using DINO features as a regularization (Kerr et al., 2023; Zuo et al., 2024), while alleviating the computational overhead associated with joint gradient-based optimization of CLIP and DINO features.
Acknowledgments
This project was supported by ANR 3IA MIAI@Grenoble Alpes (ANR-19-P3IA-0003) and by ERC grant number 101087696 (APHELEIA project). This work was granted access to the HPC resources of IDRIS under the allocation [AD011013343R2] made by GENCI.
Appendix
Appendix A Using LUDVIG for downstream tasks
In this section, we describe our approach for uplifting DINOv2, SAM and CLIP models and evaluating the 3D features on two downstream tasks: segmentation and open-vocabulary object detection.
As in Sec. 3, we are given a set of 2D frames $I_{1},\dots,I_{m}$, with viewing directions $d_{1},\dots,d_{m}$ and a corresponding 3D scene obtained using the Gaussian Splatting method.
Multiple-view segmentation.
For this task, we assume that a foreground mask of the object to be segmented is provided on a reference frame taken to be the first frame $I_{1}$. We consider two types of foreground masks: either scribbles or a whole reference mask of the object, both of which define a set of foreground pixels $\mathcal{P}$. In the following subsections, we present the proposed approaches for segmentation using SAM and DINOv2 features, based on both types of foreground masks.
A.1 Multiple-view Segmentation with SAM
SAM (Kirillov et al., 2023; Ravi et al., 2024) is a powerful image segmentation model, that can generate object segmentation masks from point prompts on a single 2D image. Aggregating SAM 2D segmentation masks in 3D allows for cross-view consistency and improves single-view segmentation results. In order to leverage SAM, we propose a simple mechanism for generating SAM 2D features for each frame from a foreground mask in the reference frame.
Generating 2D query points for SAM.
The key idea is to generate point prompts on each training frame from the foreground mask provided on the reference frame. To this end, we perform an uplifting of the foreground mask (Eq. (3)) and re-project it on all frames (Eq. (4)). From the reprojected mask for viewing direction $d$, further normalized by its average value, we retain a subset $\mathcal{P}_{d}$ of pixels with values higher than a threshold $\tau$ fixed for all scenes ($\tau=0.4$ for SPIn-NeRF and $\tau=1$ for NVOS). We then predict a SAM mask based on these point prompts as described below.
Predicting a SAM segmentation mask from a set of query points.
Given a set of pixels $\mathcal{P}_{d}$ pertaining to the foreground, we compute 2D segmentation masks using SAM by randomly selecting $3$ points prompts from $\mathcal{P}_{d}$, repeating the operation $10$ times and averaging the resulting masks for each view to obtain the final 2D SAM feature maps.
Segmentation with uplifted SAM masks.
The 2D segmentation masks generated by SAM are uplifted using the aggregation scheme described in Sec. 3.2. Our final prediction is obtained by rendering the uplifted feature maps into the target frame and thresholding.
Evaluation and hyperparameter tuning.
Segmentation with 3D SAM masks requires setting a threshold for foreground/background pixel assignment, and optionally choosing one of the three masks proposed by SAM (representing different possible segmentations of the object of interest). For SPIn-NeRF, the threshold and mask prediction are chosen based on the IoU for the available reference mask. For NVOS, only reference scribbles are provided; hence, a single mask is predicted, and the segmentation threshold is fixed across all scenes for SAM, and automatically chosen using Li’s iterative Minimum Cross Entropy method (Li & Lee, 1993) for SAM 2.
A.2 Multiple-view segmentation with DINOv2
DINOv2 (Oquab et al., 2024) is a self-supervised vision model recognized for its generalization capabilities. In this work, we aggregate the patch-level representations produced by DINOv2 with registers (Darcet et al., 2024) into a high resolution and fine-grained 3D semantic representation.
Construction of 2D feature maps.
We construct the 2D feature maps using a combination of a sliding windows mechanism and dimensionality reduction of the original DINOv2 features. Specifically, we i) extract DINOv2 patch-level representations across multiple overlapping crops of the training images, ii) apply dimensionality reduction over the set of all patch embeddings, ii) upsample and aggregate the dimensionality-reduced patch embeddings to obtain pixel-level features for each image. The process is illustrated in Figure 4. This approach enhances the granularity of spatial representations by aggregating patch-level representations to form pixel-level features. To favor the first principal components, known to focus on the foreground objects (Oquab et al., 2024), the features are re-weighted by the eigenvalues of the PCA decomposition.
Segmentation with uplifted DINOv2 features.
The 2D feature maps from the $m$ training views are uplifted using Eq. (3) and the resulting 3D features are then re-projected into any viewing direction $d$ using Eq. (4) to compute rendered 2D features $(\hat{F}_{d,p})$. To obtain segmentation masks, we construct a score $P(\hat{F}_{d,p})$ for a 2D pixel $p$ to belong to the foreground, based on its corresponding rendered feature. More precisely, $P$ relies on the rendered foreground features $\mathcal{F}_{\text{ref}}:=(\hat{F}_{d_{1},p})_{p\in\mathcal{P}}$ corresponding to the foreground mask computed on the reference frame $I_{1}$. We propose two approaches for constructing $P$. The first one is a simple approach that sets $P(\hat{F}_{d,p})=\mathcal{S}_{F}(\hat{F}_{d,p},\bar{F})$ where $\bar{F}$ is the average over foreground features $\mathcal{F}_{\text{ref}}$, and $\mathcal{S}_{F}$ is defined based on the cosine similarity. The second approach is more discriminative and first trains a logistic regression model $P$ on all rendered 2D features of the reference frame, so that the foreground features $\mathcal{F}_{\text{ref}}$ are assigned a positive label. Then $P(\hat{F}_{d,p})$ gives the probability that a pixel $p$ belongs to the foreground. The final mask is then obtained by thresholding.
Experimentally, the second approach is extremely efficient when the set of foreground pixels $\mathcal{P}$ covers the whole object to segment, so that $P$ captures all relevant features. This is the case when a whole reference mask of the object is provided. When the foreground pixels $\mathcal{P}$ does not cover the whole object, as with scribbles, $P$ can be discriminative to parts of the object that are not covered by $\mathcal{P}$. Therefore, we rely on the second approach for tasks where a reference mask is provided, and use the simpler first approach when only scribbles serve as reference.
Figure 4: Sliding windows for construction of DINOv2 feature maps.
A.3 Enhancing segmentation with DINOv2 using 3D graph diffusion
DINOv2 provides generic visual features that do not explicitly include information for segmentation, unlike models such as SAM that were specifically trained for such a task. Consequently, using the 2D projections of uplifted DINOv2 features, as proposed in Sec. A.2, might fail to separate different objects that happen to have similar features while still being distinct entities. This challenge can be mitigated by incorporating 3D spatial information in which the objects are more likely to be well-separated. To this end, we propose to leverage the graph diffusion process introduced in Section 3.3. Below, we describe the initialization of the weight vector $g_{0}$ and the construction of the adjacency matrix $A$.
Initialization of the weight vector.
The initial vector of weights $g_{0}\in\mathbb{R}^{n}$ representing a coarse estimation of the contribution of each Gaussian to the segmentation mask. It is computed by uplifting the 2D foreground mask (either scribbles or a reference mask) from the reference frame into 3D using Eq. (3) and retaining only the top 10% of Gaussians with positive mask values, setting the rest to zero. The nodes for which $g_{0}$ has a positive value define a set of anchor nodes $\mathcal{M}$ that are more likely to contribute to the foreground. The resulting weight vector is a coarse estimation of how much each Gaussian contributes to a rendered 2D segmentation mask.i
Construction of the graph edges.
We define the pairwise similarity function $S_{f}$ as:
$S_{f}(f_{i},f_{j})=\exp\left(-\frac{||f_{i}-f_{j}||_{2}^{2}}{bs_{f}^{2}}\right)$ (9)
where $f_{i},f_{j}$ are the $l_{2}$-normalized DINOv2 features, $s_{f}$ is the median of pairwise $l_{2}$ distances and $b$ is a bandwidth parameter. We choose a global unary regularization term $P(f_{i})$ on each node $i$ contain diffusion to nodes with features similar to those of the foreground. More precisely, $P$ is defined using a similar approach as in Sec. A.2:
  1. 1.
    When scribblers are provided, $P(f_{i})=\mathcal{S}_{f}(f_{i},\bar{f})$ with the averaged feature $\bar{f}$ over the anchor nodes $\mathcal{M}$, and a different value for the bandwidth $b$.
  2. 2.
    When a full foreground mask is available, we train a logistic regression model on the uplifted features with positive labels for the anchor nodes’ features. The unary term is then defined as $P(f_{i})=\mathcal{L}(f_{i})^{1/b}$, with $b$ a bandwidth parameter and $\mathcal{L}(f_{i})$ is the model’s predicted probability for $f_{i}$.
The local term $S_{f}$ allows diffusing to neighbors that have similar features while the unary term prevents leakage to background nodes and allows using an arbitrary number of diffusion steps.
The matrix $A$ of graph edges is then defined based on $S_{f}$ and $P$ as in Eq. (7). For this task, we binarize $A$ with a fixed threshold (set to $10^{-5}$). After the $T$ diffusion steps, we recover the nodes $\mathcal{S}$ in $g_{T}$ with strictly positive values (i.e., those reachable after $T$ iterations). The final weight is defined as $h_{i}=P(f_{i})$ if $i\in\mathcal{S}$ and 00 otherwise. Segmentation is then performed by projecting $\mathbf{h}=(h_{i})$ into 2D and thresholding. The selection process of the threshold and bandwidth parameters is detailed below.
Evaluation and hyperparameter tuning.
Segmentation relies on three hyperparameters: the bandwidths for $S_{f}$ and $P$, and the threshold for foreground/background pixel assignment. For SPIn-NeRF, all hyperparameters are chosen based on the IoU for the available reference mask. For NVOS, only reference scribbles are provided, hence we predict a SAM mask based on the scribbles of the reference frame, and choose the hyperparameters maximizing the IoU with this SAM mask. This is consistent with a scenario where the user, here SAM, would choose hyperparameters based on visual inspection on one of the frames.
A.4 Open-vocabulary object localization
For the open-vocabulary object localization task, multi-resolution CLIP feature maps are constructed as described in Sec. 4.2 and uplifted along with DINOv2 features using Eq. (3). The refined 3D CLIP features are then evaluated on the LERF localization and segmentation tasks as described below.
A.4.1 Relevancy scores and object localization.
We consider uplifted 3D CLIP features $f$. We follow LERF (Kerr et al., 2023) and LangSplat (Qin et al., 2024) to compute alignment scores between CLIP visual features and a text query, denoted as relevancy score, and for object localization based on these relevancy scores.
Relevancy scores.
The relevancy $r_{i,q}$ between a feature $f_{i}$ and text query $\phi_{\text{q}}$ is defined as follows:
$r_{i,q}=\min_{k}\frac{\exp(T\cdot f_{i}\cdot\phi_{q})}{\exp(T\cdot f_{i}\cdot\phi_{q})+\exp(T\cdot f_{i}\cdot\phi_{cano}^{k})},$ (10)
where $T$ is a temperature parameter set to 10 by  Kerr et al. (2023) and $\phi_{\text{cano}}^{k}$ is the CLIP embedding of a predefined canonical phrase chosen from “object,” “things,” “stuff,” and “texture.” Note that Qin et al. (2024) compute the relevancy scores for 2D pixels, while we directly compute them for 3D Gaussians, allowing their manipulation in 3D.
Localization.
The 3D CLIP relevancies can be projected into 2D for a given camera pose, and used for localization and segmentation for each text query. For localization, we follow Qin et al. (2024) and choose the pixel with the highest relevancy score, following a 2D smoothing using a mean filter with kernel size $K=5$ in our work.
A.4.2 Object segmentation
Segmentation based on raw CLIP relevancies is challenging, as fully covering the object of interest without capturing other objects of a similar nature is challenging.
We first describe our process for segmenting directly based 3D relevancies. We then present two complementary approaches that allow for a more targeted segmentation: predicting 2D SAM masks by retrieving query points with high relevancies, and refining 3D relevancy scores using graph diffusion based on 3D DINOv2 features.
Segmentation from 3D relevancies.
Given a camera pose and a text query, a segmentation mask is obtained by first applying a rough thresholding over projected relevancies rescaled by their maximum value, with a fixed threshold value $\tau=0.8$, followed by automatic thresholding with Otsu’s method (Otsu et al., 1975).
Improving 2D segmentation with SAM.
For segmentation with SAM, we use the pixels with the highest relevancy scores as query points for a given camera pose and text query. More precisely, we first obtain a mask $\mathcal{M}$ by projecting and thresholding the relevancies as described above, using $\tau=0.93$. We then use the approach described in Sec. A.1, paragraph Predicting a SAM segmentation mask from a set of query points and average 20 mask predictions. We choose the top-$q$ percent of pixels as the subset of query points for SAM, where $q$ is the proportion of positive pixels in $\mathcal{M}$, hence extracting a larger subset of point prompts for larger objects. The scalar map obtained by averaging the 20 predicted masks is then automatically thresholded again using Otsu’s method (Otsu et al., 1975).
Refining 3D CLIP relevancies with graph diffusion based on DINOv2 features.
We refine 3D CLIP relevancy scores using graph diffusion based on 3D DINOv2 features ($f$), as in Sec. A.3. The diffusion process runs in parallel for all text queries. For initialization, we keep positive a very restricted set of nodes with the highest relevancy, whose weight propagate to neighboring nodes with similar DINOv2 features, progressively spanning the object of interest. The diffusion process results in a better segmentation both with and without leveraging SAM. When using SAM, the set of query pixels can have a larger coverage of the object of interest without extending to other objects.
Details on graph construction and node initialization for refining 3D CLIP relevancies.
The pairwise similarity function $S_{f}$ is defined as in Eq. 9 with a bandwidth value $b=0.5$. For each text query $\phi_{q}$, we define a unary regularization term $P_{q}$ using a logistic regression model $\mathcal{L}_{q}$ that predicts the probability that a DINOv2 feature $f_{i}$ belongs to the object corresponding to query $\phi_{q}$. The set of nodes $\mathcal{P}$ with positive labels is defined based on 3D CLIP relevancy scores for prompt $\phi_{q}$. More precisely, we rescale 3D relevancies to $[0,1]$ and apply Otsu’s method (Otsu et al., 1975) over relevancies above $0.5$. We use a regularization $C=0.001$ and equal class weighting for training the model. The unary regularization term $P_{q}$ is then defined as $P_{q}(f_{i})=\mathcal{L}_{q}(f_{i})^{1/b}$, with $b=0.025$ for segmentation with SAM, and $b=2$ otherwise. The initial weight vector $g_{0}$ is defined by applying two more iterations of Otsu’s method among nodes in $\mathcal{P}$ and setting to zero all relevancies below the given threshold. Restricting the set of initial points ensures diffusion only happens within the object of interest. Segmentation based on the resulting 3D relevancy scores is then performed as described in the previous paragraphs, using $\tau=0.01$ for segmentation with SAM and $\tau=0$ otherwise.
Appendix B Additional results
B.1 Runtime analyses
In this section, we detail our running times for feature uplifting and evaluation, conducted on a GPU A6000 ADA. Table 5 shows a breakdown of running times between feature uplifting (Up.) and generation (Gen.), graph diffusion and 2D segmentation for evaluation on LERF segmentation. The total reported times can be divided between pre-uplifting, uplifting and post-uplifting. In our experiments, the pre-processing and uplifting steps are independent from the downstream tasks (except for our foreground/background segmentation with SAM), and part of the graph-diffusion process is task-dependent. Below we detail our runtimes for every step and compare them to the literature.
Pre-uplifting: feature map generation.
The time this step takes (Gen. in Table 5) depends on the backbone model, on the number of training images and on the number of calls to the model per image. The total time ranges from a few seconds up to an hour for approaches such as LangSplat (Qin et al., 2024), that queries SAM over a grid of points on the image at various resolutions to generate full image segmentation masks. This process takes 24s/image on a GPU 6000 ADA and amounts to an average of 80 minutes for the evaluated scenes. In our experiments, the feature generation step takes from 1 to 5 minutes.
Uplifting.
For LUDVIG, uplifting time is linear in the number of images (given a 3D scene representation). Experimentally, it is also linear in the number of feature dimensions, taking 2ms per dimension for an image of size $724\times 986$. As reported in Table 5 (Up.), uplifting 100 images of dimension $40$ takes 9s on average. By contrast, gradient-based optimization requires approximately $n_{\text{steps}}$ times this duration, where the number of gradient steps $n_{\text{steps}}$ typically ranges from 3,000 to 30,000 for 3D feature distillation (Kerr et al., 2023; Qin et al., 2024; Zuo et al., 2024). Gradient-based optimization can still be very fast for low-dimensional features such as SAM masks (can take as little as a few seconds, as reported by SA3D-GS (Cen et al., 2023b)) or dimensionality-reduced features (LangSplat (Qin et al., 2024) trains an autoencoder to reduce the CLIP feature dimension from 512 to 3 and runs for 25 minutes). However, optimization becomes intractable for high-dimensional features such as CLIP and DINO; FMGS (Zuo et al., 2024) relies on an efficient multi-resolution hash embedding of the scene; however, their total training time still amounts to 1.4 hours.
Post-uplifting: graph diffusion.
After uplifting, LUDVIG performs graph diffusion using pairwise DINOv2 feature similarities for segmentation tasks. In Table 5, we divide runtimes in two categories:
  • Scene: operations performed once for the whole scene. This includes querying the Euclidean neighbors for each node, which is log-linear in the number of Gaussians. With 600,000 Gaussians as in our experiments, the step takes about 30s, and can be further optimized by using approximate nearest neighbor search algorithms (Wang et al., 2021). Defining $S_{f}$ based on DINOv2 features is also independent from the downstream task.
  • Prompt: operations that are specific to the downstream task. This includes defining the regularization $P$ (e.g. training logistic regression model(s)) and running the diffusion. The time taken depends on dimension of the diffused features (e.g. number of text queries): 1 to 2 seconds for foreground/background segmentation (a single mask) and 18 seconds on average for LERF segmentation (14 to 21 text queries).
Post-uplifting: segmentation.
Our evaluation on LERF involves 2D segmentation with SAM based on 3D relevancy scores. The runtime depends on the number of test images and on the total number of 2D text queries, as it involves one call to the SAM backbone per test image, and multiple calls to the SAM prediction head per text query, as detailed in Appendix A.4. Our total inference time per scene is of $8s$ on average, against $0.8s$ when not using SAM. In contrast, Langsplat does not rely on SAM at inference time, but relies on a computationally expensive feature map generation process, with more than 1 hour runtime.
Images (#) Text queries (#) DINOv2 (s) CLIP (s) Graph diffusion (s) 2D segmentation (s) Total
Scene Train Test Unique Total Gen. Up. Gen.+Up. Scene Prompt w/ SAM w/o SAM (mins)
Teatime 177 6 14 59 45 14 363 42 15 9 0.9 8
Waldo 187 5 18 22 44 18 371 39 19 4 0.5 8
Ramen 131 7 14 71 40 9 227 37 14 11 1 6
Figurines 299 4 21 56 58 38 811 45 22 8 0.8 16
Table 5: Runtimes for evaluation on LERF Segmentation (Kerr et al., 2023; Qin et al., 2024). The last column (Total) reports total time, which breaks down between i) feature map generation (Gen.) and uplifting (Up.) for all training images; ii) graph diffusion, divided between scene-specific (querying neighbors, defining $S_{f}$) and prompt-specific (defining $P$, running diffusion) operations for all text queries; iii) 2D segmentation with/without SAM for all text queries across test images. We also report the number of training and test images and the number of text queries across test images.
MVSeg SA3D-GS SAGA OmniSeg3D LUDVIG (Ours)
3D representation: NeRF GS GS NeRF GS
Uplifting: SAM SAM SAM DINOv2 SAM SAM2
Orchids 92.7 84.7 - 92.3 92.6 92.2 91.0
Leaves 94.9 97.2 - 96.0 96.2 96.3 96.3
Fern 94.3 96.7 - 97.5 96.3 97.0 97.0
Room 95.6 93.7 - 97.9 95.7 96.5 96.1
Horns 92.8 95.3 - 91.5 95.1 94.5 94.8
Fortress 97.7 98.1 - 97.9 97.5 98.3 98.3
Fork 87.9 87.9 - 90.4 85.0 86.8 86.7
Pinecone 93.4 91.6 - 92.1 93.2 88.8 90.7
Truck 85.2 94.8 - 96.1 95.6 94.9 93.9
Lego 74.9 92.0 - 90.8 91.1 92.7 92.9
Average 90.9 93.2 93.4 94.3 93.8 93.8 93.8
Table 6: Segmentation (IoU) on SPIn-NeRF (Mirzaei et al., 2023) with DINOv2, SAM and SAM2.
Fern Flower Fortress HornsC HornsL Leaves Orchids Trex Average
NVOS - - - - - - - - 70.1
SA3D 82.9 94.6 98.3 96.2 90.2 93.2 85.5 82.0 90.3
OmniSeg3D 82.7 95.3 98.5 97.7 95.6 92.7 84.0 87.4 91.7
SA3D-GS - - - - - - - - 92.2
SAGA - - - - - - - - 92.6
Ours-DINOv2 84.5 95.6 97.5 97.3 93.4 96.3 91.7 84.7 92.4
Ours-SAM 85.5 97.6 98.1 97.9 94.1 96.4 73.1 88.0 91.3
Ours-SAM2 84.8 97.3 98.3 97.7 93.4 96.7 73.1 89.1 91.3
Table 7: Segmentation (IoU) on NVOS (Ren et al., 2022) with DINOv2, SAM and SAM2.
Geometry only Single view Uplifting Graph diffusion
Model: Reference mask DINOv2 SAM2 DINOv2 SAM2 DINOv2
Orchids 71.3 91.5 78.4 91.5 91.0 92.6
Leaves 72.4 89.3 96.6 94.1 96.3 96.2
Fern 93.9 95.1 96.7 96.7 97.0 96.3
Room 77.4 95.4 95.6 97.3 96.1 95.7
Horns 80.7 90.9 93.0 94.2 94.8 95.1
Fortress 94.3 96.8 97.7 98.6 98.3 97.5
Fork 67.5 85.6 80.5 88.8 86.7 85.0
Pinecone 56.5 92.8 67.8 89.6 90.7 93.2
Truck 60.1 83.6 90.9 95.2 93.9 95.6
Lego 57.3 64.4 89.0 69.9 92.9 91.1
Average 73.1 88.5 88.6 91.6 93.8 93.8
Table 8: Segmentation (IoU) on SPIn-NeRF (Mirzaei et al., 2023). We compare purely geometrical reference mask uplifting and reprojection, single-view prediction, feature/mask uplifting, and graph diffusion leveraging DINOv2 or SAM2.
B.2 Per-scene foreground/background segmentation results
In this section, we present per-scene segmentation results on NVOS and SPIn-NeRF in Tables 67 and 8, along with an extended analysis of these results.
Segmentation on SPIn-NeRF.
We report our segmentation results for the SPin-NeRF dataset (Mirzaei et al., 2023) in Table 6. Our results are comparable to the state of the art while not relying on optimization-based approaches. Surprisingly, our segmentation with DINOv2 using graph diffusion also gives results on par with models leveraging SAM masks. Our lower segmentation results compared to OmniSeg’s can be partly attributed to poor Gaussian Splatting reconstruction of highly specular scenes such as the Fork, in which semi-transparent Gaussians floating over the object try to represent reflections or surface effects that are difficult to capture with standard rasterization techniques (Jiang et al., 2024).
Segmentation on NVOS.
We report our segmentation results for the NVOS dataset (Ren et al., 2022) in Table 7. Our results are comparable to those obtained by prior work. Again, DINOv2 performs surprisingly well while not having been trained on billions of labeled images like SAM. Compared to SAM, DINOv2 better captures complex objects, but sometimes also captures some background noise. This can be seen in Appendix Figure 5 with the example of Trex: while SAM misses out the end of the tail as well as the end of the ribs, DINOv2 captures the whole Trex, but also captures part of the stairs behind. Visualisations of Orchids in Appendix Figure 5 also explain the lower performance of SAM on this scene: the two orchids SAM is missing are not covered by the positive scribbles, which makes the task ambiguous.
Ablation study.
In Table 8, we compare our segmentation protocol using DINOv2 and SAM2 to multiple simpler variants. More precisely, we evaluate i) a purely geometrical variant that does not use SAM2 or DINOv2, ii) single-view segmentation in 2D based on SAM2 or DINOv2 2D predictions, iii) uplifting DINOv2 features or SAM2 masks into 3D then rendering them for segmentation, as described in Sec. A.1 and A.2, and iv) segmenting using graph diffusion over DINOv2 3D feature similarities.
The purely geometrical approach works well on the forward-facing LLFF scenes (Orchids to Fortress). In these scenes, the reference mask is accurately uplifted and reprojected as the viewing direction changes only a little between each frame. However, it fails on the 360-degree scenes (Fork, Pinecone, Truck, Lego). This points to a suboptimal 3D reconstruction of the scene, likely due to overfitting on the limited numbers of available training views (Chung et al., 2024).
The single-view variants use a similar process for constructing the features and using them for segmentation as in Sec. A.1 and A.2 but without uplifting and rendering. It improves from a purely geometrical approach and performs reasonably well on average, the foreground being well isolated from the rest of the scene. However, as illustrated in Figure 2, the semantic features are at a much lower resolution than those resulting from 3D uplifting, leading to a coarser segmentation.
3D uplifting considerably boosts results compared to single-view approaches. However, performing segmentation in 2D based on the uplifted DINOv2 features does not benefit from the 3D spatial information and typically fails on the 360-degree scenes (Pinecone, Truck and Lego) which have higher variability between frames from different views. Introducing 3D spatial information through 3D graph diffusion results in a boosted performance on these scenes.
Appendix C Additional visualizations
C.1 Segmentation tasks
Segmentation on NVOS.
Figure 5 shows our segmentation masks from SAM and DINOv2 for the three most challenging scenes of the NVOS dataset: Fern, Orchids and Trex.
(a) Reference image
(b) Ground truth mask
(c) DINOv2 mask
(d) SAM mask
Figure 5: Segmentation results on NVOS (Ren et al., 2022) with DINOv2 and SAM.
Diffusion process.
Figure 6 illustrates different steps of the diffusion process for Fern, Leaves, Flower and Trex from the NVOS (Ren et al., 2022) dataset. Starting from the reference scribbles, the diffusion rapidly spreads through the large neighboring Gaussians. Covering the entire object takes more time for complex structures such as Fern, or for masks with disconnected components such as Orchids. As illustrated in the case of Flower, the last diffusion steps allow spreading to the smaller Gaussians on the flowers’ edges, yielding a refined segmentation mask. For Trex, the parts being reached the latest are the head and tail. Their features are further away from the reference features (defined as the average feature over 3D reference scribbles), and therefore the regularization for diffusion is stronger in these regions. Overall once the object has been fully covered, the regularization is very effective at preventing leakage, which allows diffusion to run for an arbitrary number of steps.
Figure 6: Illustration of the graph diffusion process. 2D projections of i) first three PCA components of DINOv2 3D features, ii) unary regularization term (red), iii) weight vector $g_{t}$ at timesteps $t\in\{0,3,5,10,100\}$, iv) RGB segmentation obtained using a mask based on the 2D projection of $g_{100}$.
Scene editing.
Figure 7 shows comparative visualizations of scene editing with N3F (Tschernezki et al., 2022) and LUDVIG. For rendering the edited RGB image, N3F sets to zero the occupancy for all 3D points belonging to the object. For LUDVIG, we remove all Gaussians pertaining to the 3D semantic mask resulting from graph diffusion. We observe that the regions behind to segmented object are much smoother for LUDVIG than for N3F. Regions unseen from any viewpoint are black for LUDVIG (no gaussians) and result in a background partially hallucinated by NeRF for N3F.
(a) RGB image
(b) N3F (Tschernezki et al., 2022)
(c) LUDVIG
Figure 7: Scene editing. 3D segmentation, removal and rendering for LUDVIG and N3F (Tschernezki et al., 2022). For N3F, figures are sourced from (Tschernezki et al., 2022).
C.2 Visual comparisons of uplifted features
Figure 8 show a comparison of LUDVIG’s 3D DINOv2 features with learned 3D DINO features of N3D (Tschernezki et al., 2022). Their figures are taken from their work. The notable differences are a more fine-grained reconstruction of the background for the trex and horns, and overall smoother features across all scenes.
Figure 8: Comparison between LUDVIG’s uplifted DINOv2 features (bottom) and N3F’s (Tschernezki et al., 2022) learned DINO features (top). For N3F, figures are sourced from  (Tschernezki et al., 2022).
C.3 Comparison to GaussianEditor’s uplifting.
Our aggregation procedure in Eq. (5), illustrated in Figure 1, bears similarity with the one from Chen et al. (2024) for uplifting 2D binary masks to a 3D Gaussian splatting scene. In their method, uplifted masks are thresholded to create 3D binary masks that are used for semantic tracing. Specifically, they rely on rough 3D segmentation masks to selectively optimize Gaussians that are relevant for an editing task. Unlike in Eq. (3) and (5), Chen et al. (2024) propose to normalize their uplifted masks based on the total count of view/pixel pairs $(d,p)$ contributing to the mask of a Gaussian $i$, i.e. $\sum_{d,p\in\mathcal{S}_{i}}1$, without taking the rendering weight $w_{i}(d,p)$ into account. Consequently, the uplifted features tend to have larger values for large, opaque Gaussians. Figure 9 shows a qualitative comparison between 3D DINOv2 features obtained using the aggregation proposed by Chen et al. (2024) and our approach. The aggregation by Chen et al. (2024) fails to assign the right semantics to large gaussians, which is particularly visible in scenes with high specularity such as Room. This showcases the importance of defining 3D features as convex combinations of 2D pixel features.
(a) GaussianEditor
(b) LUDVIG
(c) GaussianEditor
(d) LUDVIG
Figure 9: Comparison to GaussianEditors’s uplifting. Comparison of PCA visualization of uplifted features between LUDVIG’s and GaussianEditor’s aggregation (Chen et al., 2024).
C.4 Visualization of CLIP localization and segmentation results
In this section, we present illustrations of the impact of the diffusion process (Figure 10), and comparative visualizations of localization heatmaps with LangSplat and LERF (Figure 11).
C.4.1 Impact of segmentation with SAM and DINOv2-guided graph diffusion
Figure 10 shows 2D segmentation masks colored by CLIP relevancy scores, obtained with and without leveraging SAM and/or DINOv2-guided graph diffusion for refining 3D relevancy scores.
Direct segmentation from raw 3D relevancy scores.
Isolating a specific object in the scene directly based on CLIP relevancy scores is challenging: the segmentation masks obtained by automatic thresholding include parts of other objects with similar features, like for the sheep (segmentation of the bear nose) and the spoon. The segmentation might also cover surroundings of the object of interest simply due to the low resolution of CLIP visual features, such as in the knife example.
2D segmentation with SAM.
SAM delivers a precise 2D segmentation of the object covered by points with the highest relevancy scores. However, point prompts may not span the entire object, resulting in undersegmentation, like for the sheep. In some cases, point prompts with the highest relevancy may even be located on the wrong object, resulting in an entirely wrong segmentation (e.g., the bowl segmented instead of the spoon).
Relevancy score refinement with graph diffusion based on 3D DINOv2 features.
The graph diffusion process starts with positive weights for Gaussians with the highest relevancy scores, and propagates their weight to neighbors with similar DINOv2 features. However in cases where the object of interest consists of multiple subparts (e.g. for the sheep), the final distribution of weights may be inhomogeneous and the automatic thresholding may select only one subpart. Also, if multiple close objects are to be segmented (e.g. with the knife), the final weights may cover surrounding Gaussians and the final thresholding might not clearly isolate the objects.
3D graph diffusion combined with 2D SAM segmentation.
Combining 3D graph diffusion and 2D SAM segmentation helps solving the aforementioned problems observed when using either of the two approaches individually. The diffusion process allows selecting a large set of point prompts for SAM spanning the object of interest without covering other object with similar features, resulting in an accurate segmentation.
C.4.2 Qualitative comparison of open-vocabulary objet localization.
Figure 11 illustrates open-vocabulary object localization with LERF (Kerr et al., 2023), LangSplat (Qin et al., 2024) and LUDVIG. Both LangSplat and LUDVIG correctly localize all four example objects. For queries such as the chopsticks, LangSplat’s localization is more precise, as the CLIP features are constructed by generating full image segmentation masks with SAM. This process is computationally expensive, as constructing a full segmentation mask requires querying SAM over a grid of points on the image and takes about 23s for a single image (on a GPU A6000 ADA), which amounts to an average of 80 minutes for a scene from the LERF dataset. However, it yields coherent instance-level CLIP representations, which is desirable for downstream segmentation tasks.
Graph diffusionSAMRefer to captionRGB imageRefer to captionGround truthRefer to captionCLIP featuresRefer to captionDINOv2 featuresRefer to captionNo diffusion, no SAMRefer to captionDiffusion, no SAMRefer to captionNo diffusion, SAMRefer to captionDiffusion, SAM
Graph diffusionSAMGraph diffusionSAMRefer to captionRGB imageRefer to captionGround truthRefer to captionCLIP featuresRefer to captionDINOv2 featuresRefer to captionNo diffusion, no SAMRefer to captionDiffusion, no SAMRefer to captionNo diffusion, SAMRefer to captionDiffusion, SAM
Graph diffusionSAMGraph diffusionSAMGraph diffusionSAMRefer to captionRGB imageRefer to captionGround truthRefer to captionCLIP featuresRefer to captionDINOv2 featuresRefer to captionNo diffusion, no SAMRefer to captionDiffusion, no SAMRefer to captionNo diffusion, SAMRefer to captionDiffusion, SAM
Figure 10: Open-vocabulary object segmentation with and without using 3D graph diffusion (blue) and/or 2D SAM segmentation (green). Projections of 3D CLIP and DINOv2 features colored by three main PCA components and 2D segmentation masks colored by relevancy scores.
Figure 11: Qualitative comparisons of open-vocabulary 3D object localization on the LERF dataset. The red points are the model predictions and the black dashed bounding boxes denote the annotations. This figure is sourced and adapted from LangSplat’s website (https://langsplat.github.io/), licensed under a Creative Commons Attribution-ShareAlike 4.0 International License.